Divide 6840 between x,y and z in such a way that x has twice as much as y, who has half as much as z
Answer by ikleyn(53765)

.

If y has W dollars, then x has 2W dollars and z has 2W dollars.


The total then is  x + y + z = 2W + W + 2W = 5W = 6840 dollars.


Hence,  W = 6840/5 = 1368 dollars.


Thus y has  1368 dollars.


x and z have 2W = 2*1368 dollars EACH.